4. The managers of Bay View Bank asks for a performance/risk analysis, and asks you to answer the following questions. The Bay View Banks balance sheet is as follows: Assets: Securities Long-term Loans 5% rate $200 million S 800 million $1000 million Ave. Duration 2 vear 5 years 2% rate Total Assets Liabilities & Equity Short-term Deposits.% rate Certificates of eposit 2% rate $500 million 400 million S900 million 1 year 2 vear Total Liabilities Equity Total Liab Equity 1000 million What is the banks expected net interest income S (NII) and expected net interest margin (NIM)? [Hint: NII- Sum (Each asset x its rate) and NIM-N11 / Earning Total Assets (excludes cash) NII (Ss) a. Sum (Each liability x its rate)] NIM % b. If the bank has the NIM % that you calculated above, a PLL% of 0.50%, and a Burden % of 1.00%, what is the banks operating ROA before taxes (NIM-Burden% PLL%)? Operating ROA (OROA) c, what is the equity multiplier (EM) for the bank? (hint EM = total assets/equity) EM c. Using this equity multiplier, what is the banks Operating ROE? (hint ROE-OROA x M·Operating ROE What is the banks 1-year income (funding) gap (Rate Sensitive Assets (RSA) for 1 year - Rate Sensitive Liabilities (RSL) for 1 year? Funding Gap a. e. Given this funding gap if rates go up by 1%, what is the expected change in the banks NII $? [Hint: Change N11 $ Funding Gap x Change Rate Expected Change in NII e. What is the Banks Duration gap (D-Gap)? D-GAP = Duration of Assets _ {[Total Liabs-Total Assets] x Duration Labs.] Hint: Duration of Assets- Sum {[Each type of asset /Total Assets] x its Duration) Duration of Liabilities - Sum {[Each type of Liability/Total Liabs.] x its Duration) Duration of Assets Duration of Liabilities Duration Gap what is the expected % change in the value of equity with a rise in rates of 1%? Expected Change in Value of Equity- D-GAP x ilChg rate / (1+ Ave loan rate)] (Use A40% as the average loan rate). Expected %Ch in the Value of the Banks Equity
Answer #1

NII = sum(each asset * rate) - sum(each liability * rate) = (200*2% + 800*5%) - (500*1%+400*2%) = 44 - 13 = 31Million

NIM% = 31/1000 = 3.1%

ORAO = 3.1%-1% - 0.5% = 1.6%

EM = total assets/total equity = 1000/100 = 10

ROE = OROA * EM = 16%